# Scaling Low-Resource MT via Synthetic Data Generation with LLMs

### v1

@@ Abstract
We investigate the potential of LLM-generated synthetic data for improving low-resource machine translation (MT). Focusing on seven diverse target languages, we construct a document-level synthetic corpus from English Europarl, and extend it via pivoting to 147 additional language pairs. Automatic and human evaluation confirm its high overall quality. We study its practical application by (i) identifying effective training regimes, (ii) comparing our data with the HPLT dataset, and (iii) testing its utility beyond English-centric MT. Finally, we introduce SynOPUS, a public repository for synthetic parallel datasets. Our findings show that LLM-generated synthetic data, even when noisy, can substantially improve MT performance for low-resource languages.
@@
[…]
@@ 1 Introduction
This raises the question: Can an MT system trained on LLM-generated data benefit truly low-resource language pairs? To date, there is little to no systematic investigation of (a) generating large-scale synthetic data using LLMs for low-resource languages, (b) evaluating its intrinsic quality, (c) quantifying its downstream impact when training or fine-tuning modern MT systems. This paper provides such a systematic investigation. We make the following main contributions:
@@
    4. 4)
      Fine-tuned models that are 10-20 times smaller than state-of-the-art multilingual models perform similar or better than their large counterparts.
  • We extend our dataset into a multi-parallel corpus via pivoting and, as a case study, we demonstrate that Finnish-Somali translation improves by +14.78 ChrF and +21.64 ChrF when fine-tuning OPUS-MT.
  • To promote reproducibility and future research, we introduce SynOPUS, a public repository of synthetic parallel corpora. We also publicly release our datasets222 https://opus.nlpl.eu/synthetic/Europarl.php The data is subject to the terms and conditions defined by the usage policies of OpenAI. and code. [^3] https://github.com/Helsinki-NLP/low-res-lmt
@@ 2 Related Work
Low-resource MT targets language pairs with little to no parallel data available Haddow et al. (2022). To mitigate the data scarcity problems, two main lines of research have emerged: (a) transfer learning Zoph et al. (2016) and multilingual training Johnson et al. (2017) and (b) data augmentation Xia et al. (2019). First, transfer learning involves using a model trained on a high-resource language as a starting point for training the low-resource language, while multilingual training proposes to train jointly on multiple language pairs to compensate for the lack of text in a specific language. Second, data augmentation proposes to generate synthetic samples to train on, by perturbing, translating or otherwise modifying existing sentences Fadaee et al. (2017). Below, we focus on data augmentation and recent work that uses LLMs to generate such data.
@@
The most popular approach for low-resource languages is back-translation, which involves translating the monolingual target-language data into the source language Ko et al. (2021); Khenglawt et al. (2024). The reverse process, forward translation of source-side monolingual sentences, has also been explored, and while less common in MT, proved valuable for LLM pretraining. For example, Wang et al. (2025) used NLLB to forward-translate monolingual corpora in nine languages and demonstrated its value for LLM pretraining. This process also relates closely to sequence-level knowledge distillation Kim and Rush (2016), where compressing a large model involves training a small student model on synthetic data constructed by forward-translating using the teacher model Gordon and Duh (2019).
@@
Despite the above advances, there is still no systematic study that produces a fully synthetic multi-parallel corpus with state-of-the-art LLMs for low-resource languages and evaluates that corpus both intrinsically and on downstream MT. We close this gap by extending Europarl, a multilingual resource with alignments across all the official EU languages, into seven low-resource languages and evaluating its quality and usefulness.
@@ 3 Dataset Construction
Table 1: Statistics of the different preprocessing steps described in Section 3.3 for each language.
Our goal is to study real-world cases instead of selecting common language pairs in an artificially constructed low-resource scenario. Therefore, we conduct a preliminary experiment to help us select the languages to prioritize (Section 3.1). We then forward-translate the English Europarl corpus (Section 3.2), and, in a final post-processing step, we filter out noise to ensure high-quality translations (Section 3.3). Finally, we expand our dataset via pivoting to all existing languages of Europarl (Section 3.4).
@@ 3.1 Language Selection
We start by selecting a small set of low-resource languages for which GPT-4o can produce usable translations. To do so, we begin with a list of 204 European minority languages [^4] The list is derived from https://en.wikipedia.org/wiki/Regional_and_minority_languages_in_Europe. and retain only the 39 languages that are supported by the FLORES-200 benchmark Goyal et al. (2022). For each of the 39 languages, we prompt GPT-4o to produce translations of (i) 100 random samples from FLORES-200 and (ii) 20 five-sentence chunks to simulate paragraph-level translation. We explicitly specify the script of the target language in the prompt [^5] Our initial experiments suggested that GPT-4o occasionally produces translations using a script different from that used in the FLORES dataset. This issue was the most prominent in Serbian, which uses both Cyrillic and Latin scripts..
To contextualize GPT-4o’s performance against existing well-performing translation models, we translate the same datasets with EMMA (Ji et al., 2024), using both zero-shot and 3-shot settings, by selecting 3 unused examples from the Flores-200 dataset. Additionally, we compare the results to the best available OPUS-MT model (Tiedemann et al., 2024) per language, selected from the OPUS-MT Dashboard Tiedemann and De Gibert (2023). We compare the performance of the three systems using ChrF Popović (2015). Table 6 presents the results of the pilot evaluation. We proceed to select seven languages that are of special interest to us: Basque (eu), Scottish Gaelic (gd), Icelandic ((is), Georgian (ka), Macedonian (mk), Somali (so), and Ukrainian (uk).
@@ 3.2 Synthetic Data Generation
We use the English Europarl [^6] To the best of our knowledge, the Europarl corpus is not subject to any copyright restrictions. Koehn (2005) as the source for generating the synthetic dataset. Europarl, which is derived from the proceedings of the European Parliament, offers well-defined document boundaries and is multi-parallel across 21 European languages. We leverage the metadata within the Europarl corpus to segment the data into paragraphs in a way that each generated translation can be matched back to its exact English source, preserving the multi-parallel structure. Paragraphs are sent in bulk to the OpenAI’s Batch API [^7] https://platform.openai.com/docs/guides/batch. We instruct the model to generate translations for source-target language pairs using the following prompt:
@@
For instance, in the English-Ukrainian direction, we set the target language (TARGET) to Ukrainian, and the script information (SCRIPT) to Cyrl.
@@ 3.3 Data Post-Processing
Because of the inherent noise in the translation process, and because of their tendencies to produce hallucinations, the LLMs may make errors in translation. To filter such cases, we apply the HeLI-OTS language identifier (Jauhiainen et al., 2022) to every generated segment after sentence segmentation, and discard those that are not classified correctly. On average, this step removes only about 0.45 % of sentences in each language.
Lastly, we align the cleaned target sentences with their English counterparts using the Yasa alignment tool (Lamraoui and Langlais, 2013), while preserving document boundaries so each sentence can be matched to its document and sentence identifiers. The resulting corpus contains 2–2.3 million aligned sentences. The data statistics are presented in Table 1.
@@
3.4 MultiEuroparl: a Multi-Parallel Document-Level Corpus
An important design decision in our experiments was the focus on an inherently multi-parallel dataset. Since all languages are aligned through English, we can use it as a pivot to project synthetic translations onto existing alignments.
@@ 3.4 MultiEuroparl: a Multi-Parallel Document-Level Corpus
Using the procedures above, we are able to create 147 new language pairs added to the original Europarl corpus, while keeping document information. All of them are now available as training data for non-English-centric MT, a valuable source that comes for "free" due to the multilinguality and metadata of the source data. We study the usefulness of this data in Section 5.4.
@@ 4.2 Human Evaluation
To further assess the quality of our synthetic dataset, we conduct a human evaluation for five languages [^9] We were, unfortunately, unable to find annotators for Scottish Gaelic and Somali for this submission. . For each of language pair, we randomly sample 100 sentences and ask native speakers to evaluate them. Each language pair was scored by 1–3 annotators. Following the Direct Assessment (DA) protocol Graham et al. (2013) from the WMT2017 Shared Task Bojar et al. (2017), annotators were shown the source sentence and its translation, and were asked to assign a score on a 0–100 scale, using the guidelines provided (Appendix B). All ratings were collected through a custom web interface built with Gradio Abid et al. (2019).
We measure Inter-Annotator Agreement (IAA) using Krippendorff’s alpha (interval level) and compute the z-scored version by normalizing each annotator’s scores to account for differences in scoring behavior. Table 2 shows the details of the human evaluation. The IAA indicates moderate consistency among annotators.
@@
Correlation scores:
We report Spearman’s rank correlation ($\rho_{s}$) between Bicleaner-AI and CometKiwi scores, and human judgments in Table 2. Overall, the correlations with BicleanerAI are weak across most language pairs ($\rho_{s}<0.26$), suggesting limited alignment. Georgian shows a relatively higher correlation ($\rho_{s}=0.39$), likely due to the greater variance in human scores for this language. We observe that Bicleaner-AI tends to assign lower scores to samples that received very high ratings from human annotators, indicating a potential underestimation of high-quality translations. In contrast, correlations are consistently higher for CometKiwi. This is expected, as CometKiwi is designed to evaluate translation quality (a task more closely aligned with human judgments).
[…]
@@
Overall, we can conclude that the generated dataset is of high quality, with both automatic and human metrics indicating that most sentence pairs are of good to excellent quality, particularly for Ukrainian and Macedonian. Lower performance is observed for Georgian, Somali, and Scottish Gaelic. This is coherent with the GPT pilot evaluation that we conducted on FLORES-200, as GPT performs the best in terms of translation accuracy on Ukrainian and Macedonian, and worse for the rest of the languages (see Table 6).
Model Language Pair ##\## Params
@@ 5 Leveraging our Synthetic Data for MT
We evaluate the quality of our synthetic data by analyzing model performance both before and after fine-tuning across multiple architectures (Sections 5.1 and 5.2), serving as a proxy for data quality. Furthermore, we compare our dataset to a web-crawled SOTA corpus (Section 5.3) and study the usefulness of MultiEuroparl (Section 5.4).
@@ 5.1 Experimental Setup
We focus on the translation direction from English into the low-resource language, as this is typically the more challenging scenario. For all experiments we use the synthetic data as training set, and the FLORES-200 Goyal et al. (2022) development and test sets for model selection and evaluation, respectively.
@@
Since the languages that we are dealing with are not linguistically similar, we train individual models for each target language. We experiment using the following models (more details are provided in Appendix C):
@@
  • NLLB-200-distilled-1.3B: the distilled 1.3B parameter NLLB-200 model (Meta AI, 2022). For the fine-tuned version, we used DeepSpeed Rasley et al. (2020).
  • LLaMA-3.2-3B-Instruct: the 3B parameter LLaMA-3.2 Instruct model (Grattafiori et al., 2024). For the fine-tuned version, we adapt LoRA (Hu et al., 2022) using Unsloth (Daniel Han and team, 2023).
@@
We evaluate all models before and after fine-tuning. We report ChrF (Popović, 2015) as our main automatic metric, as it has been the standard metric for low-resource MT and it is shown to correlate more closely with human judgments than BLEU (Papineni et al., 2002). We report COMET (Rei et al., 2020) in Appendix E.
@@ 5.2 Overall Results and Analysis
The 60M parameter baseline, trained exclusively on our dataset, surpasses the out-of-the-box performance of billion-parameter models like NLLB and LLaMA for Scottish Gaelic and Georgian, while nearly matching them for Somali and Basque. This shows that our corpus is rich enough to train functional MT systems without any external pretraining or multilingual transfer.
@@
Based on the results from the previous section, we observe high quality for Ukranian and Macedonian, medium quality for Basque and Icelandic, and noticeably lower quality for the rest.Yet, noisy does not mean useless. In fact, Table 3 shows that the languages with the noisiest synthetic corpora also result in the largest downstream gains. When the alternative is no data at all, quantity is better than quality. However, for mid-resource languages such as Macedonian and Ukrainian, where cleaner text is already available, additional synthetic data offers diminishing returns and mainly hurts the performance of these systems. The lower the resource level, the more tolerant MT training is to noise.
@@
LLaMA initially struggles (21-26 ChrF), reflecting its lack of inherent translation capability for low-resource languages. While adapter training yields substantial improvements (+19-29 ChrF), the model still underperforms smaller, translation-specific models. This indicates that while synthetic data allows for adaptation, it cannot fully compensate for mismatches between the model architecture or pretraining objective and the translation task itself In LLaMA’s case, the 3B parameter scale appears unnecessarily large for this specific MT task, and leads to unnecessarily large fine-tuning times.
@@
    Fine-tune MT multilingual models. NLLB-200 benefits consistently across all language pairs. Our findings are consistent with previous research that finds that fine-tuning NLLB is among the best approaches for low-resource MT Scalvini et al. (2025); Tapo et al. (2025); De Gibert et al. (2025).

### v2

@@ Abstract
We investigate the potential of LLM-generated synthetic data for improving low-resource Machine Translation (MT). Focusing on seven diverse target languages, we construct a document-level synthetic corpus from English Europarl, and extend it via pivoting to 147 additional language pairs. Automatic and human evaluation confirm its overall high quality. We study its practical application by (i) identifying effective training regimes, (ii) comparing our data with the HPLT dataset, (iii) studying the effect of varying training data size, and (iiii) testing its utility beyond English-centric MT. Finally, we introduce SynOPUS, a public repository for synthetic parallel datasets. Our findings show that LLM-generated synthetic data, even when noisy, can substantially improve MT performance for low-resource languages.
@@
Ona de Gibert[Uncaptioned image]   Joseph Attieh[Uncaptioned image]   Teemu Vahtola[Uncaptioned image]   Mikko Aulamo[Uncaptioned image]   Zihao Li[Uncaptioned image] Raúl Vázquez[Uncaptioned image]Tiancheng Hu[Uncaptioned image]Jörg Tiedemann[Uncaptioned image] [Uncaptioned image] University of Helsinki  [Uncaptioned image] University of Cambridge {first.last}@helsinki.fi, th656@cam.ac.uk
@@ 1 Introduction
This raises the question: Can an MT system trained on LLM-generated data benefit truly low-resource language pairs? To date, there is little to no systematic investigation of (a) generating large-scale synthetic data using LLMs for low-resource languages, (b) evaluating its intrinsic quality, (c) quantifying its downstream impact when training or fine-tuning modern MT systems. This paper provides such a systematic investigation. We make the following contributions:
@@
    4. (4)
      Fine-tuned models that are 10-20 times smaller than SOTA models perform similar or better than their large counterparts.
  • We study the effect of training data size to investigate the scalability of our approach.
  • We extend our dataset into a multi-way parallel corpus via pivoting. As a case study, we demonstrate that Finnish-Somali translation improves by +14.78 ChrF and +21.64 ChrF when fine-tuning OPUS-MT.
  • To promote reproducibility and future research, we introduce SynOPUS, a public repository of synthetic parallel corpora. We also publicly release our dataset, which we publish under a new version of Europarl (v8syn), [^2] https://opus.nlpl.eu/synthetic/Europarl.php The data is subject to the terms and conditions defined by the usage policies of OpenAI. code, [^3] https://github.com/Helsinki-NLP/low-res-lmt and baseline models. [^4] Helsinki-NLP/scaling-low-res-mt-via-synthetic-data-generation-with-llms
@@ 2 Related Work
Low-resource MT targets language pairs with little to no parallel data available Haddow et al. (2022). To mitigate the data scarcity problems, two main lines of research have emerged: (a) transfer learning Zoph et al. (2016) and multilingual training Johnson et al. (2017), and (b) data augmentation Xia et al. (2019). First, transfer learning involves using a model trained on a high-resource language as a starting point for training the low-resource language, while multilingual training proposes to train jointly on multiple language pairs to compensate for the lack of text in a specific language. Second, data augmentation proposes to generate synthetic samples to train on, by perturbing, translating or otherwise modifying existing sentences Fadaee et al. (2017). Below, we focus on data augmentation and recent work that uses LLMs to generate such data.
@@
The most popular approach for low-resource languages is back-translation, which involves translating the monolingual target-language data into the source language Ko et al. (2021); Khenglawt et al. (2024). The reverse process, forward translation of source-side monolingual sentences, has also been explored, and while less common in MT, proved valuable for LLM pretraining. For example, Wang et al. (2025) used NLLB to forward‐translate monolingual corpora in nine languages and demonstrated its value for LLM pretraining. This process also relates closely to sequence-level knowledge distillation Kim and Rush (2016), where compressing a large model involves training a small student model on synthetic data constructed by forward-translating it with the teacher model Gordon and Duh (2019).
@@
Despite the above advances, there is still no systematic study that produces a fully synthetic multi-way parallel corpus with SOTA LLMs for low-resource languages and evaluates that corpus both intrinsically and on downstream MT. We close this gap by extending Europarl, a multilingual resource with alignments across all the official EU languages, into seven low-resource languages and evaluating its quality and usefulness.
@@ 3 Dataset Construction
Table 1: Statistics of the different post-processing steps described in Section 3.3 for each language.
Our goal is to study real-world cases instead of selecting common language pairs in an artificially constructed low-resource scenario. We conduct a preliminary experiment to help us select the languages to prioritize (Section 3.1). We then forward-translate the English Europarl corpus (Section 3.2), and, in a final post-processing step, we filter out noise to ensure high-quality translations (Section 3.3). Finally, we expand our dataset via pivoting to all languages of Europarl (Section 3.4).
@@ 3.1 Language Selection
We start by selecting a small set of low-resource languages for which GPT-4o can produce usable translations. To do so, we begin with a list of 204 European minority languages [^5] The list is derived from https://en.wikipedia.org/wiki/Regional_and_minority_languages_in_Europe. and retain only the 39 languages that are supported by the FLORES+ benchmark Goyal et al. (2022). For each of the 39 languages, we prompt GPT-4o to produce translations of (i) 100 random samples from the FLORES+ development set and (ii) 20 five-sentence chunks to simulate paragraph-level translation. We specify the script of the target language in the prompt. [^6] Our initial experiments suggested that GPT-4o occasionally produces translations using a script different from that used in the FLORES+ dataset. This issue was the most prominent in Serbian, which uses both Cyrillic and Latin scripts.
To contextualize GPT-4o’s performance against existing well-performing translation models, we translate the same datasets with EMMA-500 (Ji et al., 2024), using both zero-shot and 3-shot settings, by selecting 3 unused examples from FLORES+. Additionally, we compare the results to the best available OPUS-MT model (Tiedemann et al., 2024) per language, selected from the OPUS-MT Dashboard Tiedemann and De Gibert (2023). We compare the performance of the three systems using ChrF Popović (2015).777nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no| version:2.5.1 Appendix A (Table 6) presents the results of the pilot evaluation. We proceed to select seven languages: Basque (eu), Scottish Gaelic (gd), Icelandic (is), Georgian (ka), Macedonian (mk), Somali (so), and Ukrainian (uk). We select these languages based on the linguistic diversity, low-resource coverage, model performance, and our practical interest.
@@ 3.2 Synthetic Data Generation
We use the English Europarl [^8] To the best of our knowledge, the Europarl corpus is not subject to any copyright restrictions. Koehn (2005) as the source for generating the synthetic dataset. Europarl, which is derived from the proceedings of the European Parliament, offers well-defined document boundaries and is multi-way parallel across 21 European languages. We leverage the metadata within the Europarl corpus to segment the data into paragraphs in a way that each generated translation can be matched back to its exact English source, preserving the multi-way parallel structure. Paragraphs are sent in bulk to the OpenAI’s Batch API [^9] https://platform.openai.com/docs/guides/batch. We instruct the model to generate translations for source-target language pairs using the following prompt:
@@
For instance, in the English-Ukrainian direction, we set the target language (TARGET) to Ukrainian, and the script information (SCRIPT) to Cyrl. We use script identifiers from the FLORES+ language codes, such as ukr_Cyrl.
@@ 3.3 Data Post-Processing
Because of the inherent noise in the translation process, and because of their tendencies to produce hallucinations, the LLMs may make errors in translation. To filter such cases, we apply language identification using heliport [^10] https://pypi.org/project/heliport/, which is based on the HeLI-OTS language identification models (Jauhiainen et al., 2022), to every generated segment after sentence segmentation, and discard those that are not classified correctly. On average, this step removes only about 0.45 % of sentences in each language.
Lastly, we align the cleaned target sentences with their English counterparts using the Yasa alignment tool (Lamraoui and Langlais, 2013), while preserving document boundaries so each sentence can be matched to its document and sentence identifiers. The resulting corpus contains 2–2.3 million aligned sentences per language pair. The data statistics are presented in Table 1.
@@
3.4 MultiEuroparl: a Multi-Way Parallel Document-Level Corpus
An important design decision in our experiments was the focus on an inherently multi-way parallel dataset. Since all languages are aligned through English, we can use it as a pivot to project synthetic translations onto existing alignments.
@@ 3.4 MultiEuroparl: a Multi-Way Parallel Document-Level Corpus
Using the procedures above, we are able to create 147 new language pairs added to the original Europarl corpus, while keeping document information. Europarl has multi-way parallel corpora originally available in 21 languages. We add 7 new languages to the existing 21, yielding additional $21\times 7=147$ language pairs when pairing each new language with all the existing ones. All of the language pairs are now available as training data for non-English-centric MT, a valuable source that comes for "free" due to the multilinguality and metadata of the source data. We study the usefulness of this data in Section 5.5.
@@ 4.2 Human Evaluation
To further assess the quality of our synthetic dataset, we conduct a human evaluation for five languages. [^12] We were, unfortunately, unable to find annotators for Scottish Gaelic and Somali. For each of language pair, we randomly sample 100 sentences and ask native speakers to evaluate them. Each language pair was scored by 1–3 annotators. Following the Direct Assessment (DA) protocol Graham et al. (2013) from the WMT2017 Shared Task Bojar et al. (2017). Annotators were shown the source sentence and its translation, and were asked to assign a score on a 0–100 scale, using the guidelines provided (see Appendix B). All ratings were collected through a custom web interface built with Gradio Abid et al. (2019).
We measure Inter-Annotator Agreement (IAA) using Krippendorff’s alpha (interval level) and compute the z-scored version by normalizing each annotator’s scores to account for differences in scoring behavior. Table 2 shows the results of the IAA, which indicates moderate consistency among annotators.
@@
Correlation scores
We report Spearman’s rank correlation ($\rho_{s}$) between Bicleaner-AI and COMETKiwi scores, and human judgments in Table 2. In general, the correlations with BicleanerAI are weak across most language pairs, suggesting limited alignment. Georgian shows a relatively higher correlation ($\rho_{s}=0.39$), likely due to the greater variance in human scores for this language. We observe that Bicleaner-AI tends to assign lower scores to samples that received very high ratings from human annotators, indicating a potential underestimation of high-quality translations. In contrast, correlations are consistently higher for COMETKiwi. This is expected, as COMETKiwi is designed to evaluate translation quality (a task more closely aligned with human judgments).
Pair Ann. Count z-IAA $\rho_{s}\text{(BicleanerAI)}$ $\rho_{s}$ (COMETKiwi)
@@
Overall, we can conclude that the generated dataset is of fairly good quality, with both automatic and human metrics indicating that most sentence pairs are of good to excellent quality, particularly for Ukrainian and Macedonian. Lower performance is observed for Georgian, Somali, and Scottish Gaelic. This is coherent with the GPT-4o pilot evaluation that we conducted on FLORES+, as GPT-4o performs the best in terms of translation accuracy on Ukrainian and Macedonian, and worse for the rest of the languages (see Appendix A, Table 6).
Model Language Pair #\# Params
@@ 5 Leveraging our Synthetic Data for MT
We evaluate the quality of our synthetic data by analyzing model performance both before and after fine-tuning across multiple architectures (Sections 5.1 and 5.2), serving as a proxy for data quality. Furthermore, we compare our dataset to a web-crawled SOTA corpus (Section 5.3), investigate the effect of variable training data size (Section 5.4) and study the usefulness of MultiEuroparl (Section 5.5).
@@ 5.1 Experimental Setup
We focus on the translation direction from English into the low-resource language, as this is typically the more challenging scenario. For all experiments we use the synthetic data as training set, and the FLORES+ Goyal et al. (2022) development and test sets for model selection and evaluation, respectively.
@@
Since the languages under consideration are not linguistically similar, we train individual bilingual models for each target language and leave multilingual studies for future work. We experiment using the following models (more details are provided in Appendix C):
@@
  • NLLB-200-distilled-1.3B: the distilled 1.3B parameter NLLB-200 model (Meta AI, 2022). For fine-tuning NLLB, we used DeepSpeed Rasley et al. (2020).
  • Llama-3.2-3B-Instruct: the 3B parameter Llama-3.2 Instruct model (Dubey et al., 2024). For the fine-tuned version, we adapt LoRA (Hu et al., 2022) using Unsloth (Han et al., 2023).
@@
We evaluate all models before and after fine-tuning. We report ChrF (Popović, 2015) as our main automatic metric, as it has been the standard metric for low-resource MT and it is shown to correlate more closely with human judgments than BLEU (Papineni et al., 2002). We report COMET [^13] We use the Unbabel/wmt22-comet-da model. (Rei et al., 2020) for all our experiments in Appendix E.
We also evaluate GPT-4o on the full FLORES+ test set (in the pilot evaluation in Appendix A, we used only 100 samples from the development set) and include it as a reference in our results.
@@ 5.2 Overall Results and Analysis
The 60M parameter baseline, trained exclusively on our dataset, surpasses the out-of-the-box performance of billion-parameter models like NLLB and Llama for Basque, Scottish Gaelic, Icelandic and Georgian, while nearly matching them for Somali. This shows that our corpus is rich enough to train functional MT systems without any external pretraining or multilingual transfer.
@@
Based on the results from the previous section, we observe high quality for Ukranian and Macedonian, medium quality for Basque and Icelandic, and noticeably lower quality for the rest. Yet, noisy does not mean useless. In fact, Table 3 shows that the languages with the noisiest synthetic corpora also result in the largest downstream gains (Scottish Gaelic, Georgian and Somali). When the alternative is no data at all, quantity is better than quality. However, for mid-resource languages such as Macedonian and Ukrainian, cleaner text is already available beforehand and multilingual pretrained models benefit from large quantities of data from closely related languages. Therefore, additional synthetic data offers diminishing returns and mainly hurts the performance of these systems. The lower the resource level, the more tolerant MT training is to noise.
@@
Llama initially struggles (13-30 ChrF), reflecting its lack of inherent translation capability for low-resource languages. While adapter training yields substantial improvements (+11-29 ChrF), the model still underperforms compared to smaller, translation-specific models. This indicates that while synthetic data allows for adaptation, it cannot fully compensate for mismatches between the pretraining objective and the translation task itself. In Llama’s case, the 3B parameter scale appears unnecessarily large for this specific MT task, and leads to unnecessarily large fine-tuning times.
Competitiveness with GPT-4o
GPT-4o [^14] Since GPT-4o’s training data is not public, FLORES+ may be included, making evaluation unfair. is the best performing system for almost all language directions, however fine-tuned models like OPUS-MT-ft and NLLB-ft still offer competitive results, despite being much smaller. OPUS-MT-ft is far behind by 1-2 ChrF points in most languages and NLLB-ft even outperforms GPT-4o for Georgian. This indicates that while GPT-4o is a challenging system to beat, smaller and more efficient models can achieve comparable results with far fewer parameters. These models close the performance gap using only our moderately-sized synthetic corpus ($\approx$2M sentences), a mere fraction of the vast data required to train a frontier model like GPT-4o. This underscores a key finding: a targeted strategy of generating high-quality data provides a powerful and practical pathway to SOTA performance. Our approach significantly lowers the computational and financial barriers, making high-quality MT for low-resource languages much more accessible.
@@
    Fine-tune MT multilingual models. NLLB-200 benefits consistently across all language directions. Our findings are consistent with previous research that finds that fine-tuning NLLB is among the best approaches for low-resource MT Iyer et al. (2024); Zhu et al. (2024); Scalvini et al. (2025); Tapo et al. (2025); de Gibert et al. (2025).